OpenStudy (adamk):

The point (7, 3pi/4) in polar coordinates is how far from the x-axis? I just do 7sin(135) right? I get 4.9.

OpenStudy (anonymous):

Yup that would be right. Although you might want a more specific answer... \[\sin 135=\frac{ 1 }{ \sqrt2 }\] So \[7 \times \frac{ 1 }{ \sqrt2 }=\frac{ 7 }{ \sqrt2 }\]

OpenStudy (anonymous):

You could rationalize that if you wanted